# Learning under Imitative Strategic Behavior with Unforeseeable Outcomes

### v1

@@ B.2 More practical examples fitting to our model
B.3 The option of taking no action
The comprehensive probabilistic model can be easily extended to the setting where “manipulate", “improve" and “do nothing" are all possible. Specifically, denoting the expected utility of doing nothing as $U_{N}(\theta)$. We know $U_{M}(\theta),U_{I}(\theta)$ do not change, while $U_{N}(\theta)=0$. Thus, we can derive manipulation probability $P_{M}(\theta)=Pr(U_{M}(\theta)>U_{I}(\theta)~{}and~{}U_{M}(\theta)>0)$ and $P_{I}(\theta)=Pr(U_{I}(\theta)>U_{M}(\theta)~{}and~{}U_{I}(\theta)>0)$.
With $P_{M}(\theta),P_{I}(\theta)$, we can write out the new strategic utility and study its property if necessary. However, in reality, it is more reasonable for individuals to always take an action. For instance, applicants feeling they are not qualified will at least take some measures to improve their chances to be admitted. Thus, the model in the main paper disallows “taking no action".
@@ B.4 Discussion on adjusted preferences
Utility loss from the adjusted preferences. Although adjusting preferences is a simple yet effective way to promote fairness and disincentivize manipulation, the actual utility received by the decision-maker inevitably diminishes as $k_{1}$ or $k_{2}$ changes (as the actual utility the decision-maker receives is always determined by the original function $U(\theta)$ in Eq. (2.2)). Nonetheless, such diminished utilities may still be higher than the utility under non-strategic policy $\widehat{\theta}^{*}$. This is illustrated empirically in Sec. 5 and Appendix C.
Adjusted preference as a regularizer to promote fairness. We have shown that adjusting weights $k_{1},k_{2},k_{3}$ in learning objective (Eq. (5)) can control the individual behavior and algorithmic fairness. Indeed, we can view this adjustment mechanism as a regularization method: by adjusting weights, we are essentially changing the objective $U(\theta)$ by adding a regularizer, i.e.,
@@ B.5 Estimate Model Parameters
Robustness of results when $q,\epsilon$ are noisy. We also present an experiment to relax the assumption that the decision-maker knows $q,\epsilon$ exactly on FICO data. Instead, they only know $q+\delta$ or $\epsilon+\delta$ where $\delta$ is a Gaussian noise. We do 100 rounds of simulations and produce plots with expectation and error bars similar to Fig. 4(Fig. 7 and Table 3 show the results with noisy $q$, while Fig. 7 and Table 4 show the results with noisy $\epsilon$). The results show adjusting $k$ still works under noisy $q$ and $\epsilon$ although inconsistency exists.
Figure 6: Impact of adjusted preferences (FICO data) when there is a Gaussian noise on $q$. The noises have 00 mean, and $0.05,0.1,0.15$ standard deviation from the left two plots to the right two plots.
@@ C.1 Additional results on FICO score
Table 5: Comparison between three types of optimal thresholds for FICO data. For utility and $P_{M}$, the left value in parenthesis is for Asian, while the right is for Hispanic. The fairness metric is eqopt.
Threshold category Utility $P_{M}$ Unfairness
Non-strategic $(0.726,0.427)$ $(0.115,0.322)$ 0.089
Original strategic $(0.734,0.448)$ $(0.055,0.161)$ 0.047
Adjusted strategic $(0.726,0.434)$ $(0.023,0.070)$ 0.022
@@
Figure 10: Illustration of Thm. 4.5 and Thm. 4.6 on FICO Data (Asian vs Hispanic)
Figure 11: Illustration of (C.2)
Moreover, besides the illustration of Thm. 4.5 and Thm. 4.6 using Caucasian and African American data. We also demonstrate the same results hold for Asian and Hispanic as in Fig. 10.
Figure 12: Manipulation probability $P_{M}(\theta)$ of Asian and Hispanic
The scenario considered in Figure 12 satisfies the condition 1.(ii) in Thm. 4.5, because the original strategic optimal threshold $\theta_{s}^{*}<\theta_{max}$ for both groups. We further conduct experiments in this setting to evaluate the impacts of adjusting preferences. We consider equal opportunity (EqOpt) as the fairness metric, under which $\mathbb{E}_{X\sim P_{s}^{\mathcal{C}}}[\textbf{1}(X\geq\widehat{\theta})]=F_{X|YS}(\theta|1,s)$ and the unfairness measure can be reduced to $\big{|}F_{X|YS}(\theta|1,a)-F_{X|YS}(\theta|1,b)\big{|}$.
The results are shown in Figure 10, where dashed red and dashed blue curves are manipulation probabilities under non-strategic $\widehat{\theta}^{*}$ and strategic $\theta^{*}(k_{1})$, respectively. Solid red and solid blue curves are the actual utilities $U(\widehat{\theta}^{*})$ and $U(\theta^{*}(k_{1}))$ received by the decision-maker. The difference between the two green curves measures the unfairness between Asian and Hispanic. $k_{1}=1$ corresponds to the original decision-maker while others when $k_{1}>1$ indicate the decision-maker with adjusted preferences. Results show that compared to the non-strategic $\widehat{\theta}^{*}$, the strategic $\theta^{*}$, by taking into account strategic behavior disincentives the strategic manipulation. When condition 1(ii) in Thm. 4.5 is satisfied, increasing $k_{1}$ can disincentivize manipulation (i.e., $P_{M}$ decreases) while improving fairness. These validate Thm. 4.5 and 4.6.
In Table 5, We summarize the comparison between non-strategic $\widehat{\theta}^{*}$, original strategic ${\theta}^{*}$, and adjusted strategic ${\theta}^{*}(k_{1})$ (when $k_{1,a}=k_{1,h}=1.5$). It shows that decision-makers by adjusting preferences can significantly mitigate unfairness and disincentivize manipulation, with only slight decreases in utilities.
Experiments with demographic parity as new fairness metric
We also reconducted the above experiments with demographic parity (DP) as the new fairness metric. As illustrated in Sec. 4.3, $P_{s}^{\texttt{DP}}(x)=P_{X|S}(x|s)$. Similar to Fig. 4 and the bottom plot of Fig. 10, we produce Fig. 13 based on DP, which demonstrate the same patterns as the figures based on Eqopt.
@@ C.2 Results for Gaussian Data
To illustrate the complex nature under different permutations of parameters, with the pre-determined parameters in (C.2) and $\alpha=0.6$, we vary $q$ and plot both non-strategic optimal thresholds and regular strategic ones with respect to different $\epsilon$ as shown in the bottom plot of Fig. 14, where the lower graphs illustrate Thm. 4.1, i.e. the red line is always under the blue line.
Figure 14: (Non)-strategic optimal threshold
We also demonstrate the strategic utility under different combinations of $q,\epsilon$ with pre-determined parameters in (C.2) and $\alpha=0.3$ or $\alpha=0.6$. Fig. 16 and 16 suggest the complicated nature of regular strategic utility under different parameter combinations. It is possible to have 0,1 or 2 extreme points.
@@
To illustrate 4.5, we demonstrate the effects of adjusting each of $k_{1},k_{2},k_{3}$. According to Fig. 17 and 18, we can see when $k_{1}$ is large enough, the optimal strategic threshold is definitely lower than the optimal non-strategic ones. However, when $\alpha$ is small, we need larger $k_{1}$ to pull $\theta^{*}$ downward. According to Fig. 19 and 20, we can see when the population is majority qualified, adjusting $k_{2}$ is not guaranteed to shift $\theta^{*}$ upward (Fig. 20).
@@
Figure 18: Strategic optimal threshold $\theta^{*}(k_{1})$ after increasing $k_{1}$ while keeping $k_{2},k_{3}$ fixed. Left figure has $q=0.01$ and right figure has $q=0.99$, while both figures have $\alpha=0.3$
Figure 19: Strategic optimal threshold $\theta^{*}(k_{2})$ after increasing $k_{2}$ while keeping $k_{1},k_{3}$ fixed. Left figure has $q=0.01$ and right figure has $q=0.99$, while both figures have $\alpha=0.3$
@@
We first show a parameter setting satisfying condition 1.(i) in Thm. 4.5. With pre-determined parameters in (C.2) , we set $q=\epsilon=0.5$ and $\alpha_{a}=0.2,\alpha_{b}=0.25$. This matches the notation tradition in Sec. 4.3 where group $a$ is the disadvantaged group with a lower qualified percentage. Also, because $q+\epsilon\geq 1$, the setting satisfies condition 1.(i) in Thm. 4.5. We first illustrate the manipulation probability under optimal original strategic threshold $\theta_{s}^{*}$ as in Fig. 11. From Fig. 22, we can set $k_{1a}=k_{2a}=1.25$ to let the strategic utility still be larger than the one under non-strategic optimal threshold(i.e. the solid blue line is above the solid red line), while lower the cumulative density dramatically (i.e. the dotted green line) to admit more qualified individuals and disincentivize manipulation (i.e. the dashed blue line). The details of comparisons are shown in Table 8.
@@
With pre-determined parameters in (C.2) , we set $q=\epsilon=0.25$ and $\alpha_{a}=0.4,\alpha_{b}=0.6$. This matches the notation tradition in Sec. 4.3 where group $a$ is the disadvantaged group with a lower qualified percentage. We first illustrate the manipulation probability under optimal original strategic threshold $\theta_{s}^{*}$ as in Fig. 21. Fig. 21 reveals that 1.(ii) in 4.5 is satisfied because the orange and green points are both located before the extreme large point of $P_{M}(\theta)$. Thus, we could increase $k_{1s}$ to disincentivize manipulation while improving fairness as shown in Fig. 22. From Fig. 22, we can set $k_{1a}=k_{2a}=1.25$ to let the strategic utility still be larger than the one under non-strategic optimal threshold(i.e. the solid blue line is above the solid red line), while lower the cumulative density dramatically (i.e. the dotted green line) to admit more qualified individuals and disincentivize manipulation (i.e. the dashed blue line). In Table 6, We summarize the comparison between non-strategic $\widehat{\theta}^{*}$, original strategic ${\theta}^{*}$, and adjusted strategic ${\theta}^{*}(k_{1})$ (when $k_{1,c}=k_{1,aa}=1.25$). It shows that decision-makers by adjusting preferences can significantly mitigate unfairness and disincentivize manipulation, with only slight decreases in utilities.
@@
Besides, we also show one more parameter setting satisfying condition 2 in Thm. 4.5. With pre-determined parameters in (C.2) , we also set $q=\epsilon=0.2$ and $\alpha_{a}=0.3,\alpha_{b}=0.35$. This matches the notation tradition in Sec. 4.3 where group $a$ is the disadvantaged group with a lower qualified percentage. Also, based on Fig. 21, $q+\epsilon<1$ and $\alpha_{a},\alpha_{b}<0.5$, the setting satisfies condition 2 in Thm. 4.5. We first illustrates the manipulation probability under optimal original strategic threshold $\theta_{s}^{*}$ and non-strategic threshold $\widehat{\theta_{s}^{*}}$ as in Fig. 21. As shown in Fig. 22, for both groups, we demonstrate the manipulation probability for $\widehat{\theta}^{*},\theta^{*}$ and $\theta(k_{1})$ when $k_{1}$ varies, (non)-strategic utility and cumulative density conditioned on $Y=1$ (i.e. this measures the unfairness based on equal opportunity). This plot suggests we can find suitable $k_{2a}$ and $k_{2b}$ to disincentivize manipulation and promote fairness, while also making the utility higher than the one under non-strategic optimal threshold. From Fig. 22, we can set both $k_{2a}$ and $k_{2b}$ at $1.25$ to let the strategic utility still be larger than the utility under non-strategic optimal threshold (i.e. the solid blue line is above the solid red line), while keeping the cumulative density function closer (i.e. the green dotted line) to mitigate unfairness, and also disincentivize manipulation (i.e. the blue dashed line). The details of comparisons are shown in Table 8.
@@ Appendix D Derivations and Proofs
D.1 Derivations of (1)
$U_{M}(\theta)$ is the expected utility gain of an unqualified agent if choosing to manipulate: i. If the manipulation is not exposed, the probability of admission is $1-F_{X|Y}(\theta|1)$ because the manipulation leads the agents to get his/her new feature from $P(X|1)$, which happens at a probability $1-\epsilon$; ii. If the manipulation is exposed, the probability of admission is 0, which happens at a probability $\epsilon$; iii. If the agent does not manipulate, the probability of admission is $1-F_{X|Y}(\theta|0)$ because now his/her feature is from the unqualified population, and keep in mind that the agents will never know the exact values of his/her feature when he/she makes decisions; Then according to the total probability theorem, the expectation of utility gain $U_{M}(\theta)=(1-\epsilon)\cdot(1-F_{X|Y}(\theta|1))+\epsilon\cdot 0-(1-F_{X|Y}(\theta|0))-C_{M}$.
$U_{I}(\theta)$ is the expected utility gain of an unqualified agent if choosing to improve: i. If the improvement succeeds, the probability of admission is $1-F_{X|Y}(\theta|1)$ because the improvement leads the agents to get his/her new feature from $P(X|1)$, which happens at a probability $q$; ii. If the manipulation is exposed, the probability of admission is $1-F^{I}(\theta)$, which happens at a probability $1-q$; iii. If the agent does not manipulate, the probability of admission is $1-F_{X|Y}(\theta|0)$.
Then according to the total probability theorem, we can derive $U_{I}(\theta)$ as well. Finally, substitute above two terms into $P_{M}(\theta)=\Pr\left(U_{M}(\theta)>U_{I}(\theta)\right)$ and we get (1).
@@ D.1 Derivations of (1)
Assumption 2.2 ensures that $P_{C_{M}-C_{I}}>0$ when $\theta$ in its domain. Thus, we can directly take the derivative inside (1), we can get $(1-q)\cdot P^{I}(\theta)-(1-q-\epsilon)P_{X|Y}(\theta|1)$. To get its sign, we only need to consider $(1-q)-(1-q-\epsilon)\frac{P_{X|Y}(\theta|1)}{P^{I}(\theta)}$.

### v2

@@ B.2 More practical examples fitting to our model
B.3 The option of "doing nothing"
Our model implicitly includes the action of "doing nothing". Specifically, our model assumes improvement only succeeds with probability $q$; and the improvement cost differs among agents which is modeled as a random variable $C_{I}$. For agents who do nothing, we may consider them as those who take improvement action at zero costs but fail to improve. Because no cost is incurred and the improvement fails, the resulting features remain the same and the outcome is equivalent to "doing nothing".
Although it may be intuitive to model "doing nothing" as a separate action, we argue that it is more realistic in practice for individuals to always take an action. As pointed out by Horowitz et al. (2024), agents who decide to participate in a system will pay an unavoidable cost, and the cost is a part of the "improvement cost". On the contrary and as we argued before, unqualified individuals in our model are not given an explicit option of "doing nothing" with a deterministic 00 cost and 00 utility in the first place. In our framework where improvement is modeled as the imitative behavior in social learning (the third paragraph on page 2), all unqualified individuals staying in the system are influenced by the desire to be qualified and must do something to imitate the qualified profiles because they know they are unqualified. Otherwise, they will just quit the system and are not in our interest. In the college admission example, all unqualified students who do not give up applications and do not cheat are doing similar things: striving to learn more courses, take tests, and prepare for application packages. The only difference is the realization of the improvement outcomes (some students really improve, some do not). To concretely answer your question of $C_{M},C_{I}$, we slightly modify the "budget" to "utility" in footnote 3: For all agents already in the decision-making system, they already have a sufficiently large initial utility exceeding $C_{M},C_{I}$. This means if they choose to quit, they lose the initial utility.
@@ B.4 Discussion on adjusted preferences
Utility loss from the adjusted preferences. Although adjusting preferences is a simple yet effective way to promote fairness and disincentivize manipulation, the actual utility received by the decision-maker inevitably diminishes as $k_{1}$ or $k_{2}$ changes (as the actual utility the decision-maker receives is always determined by the original function $U(\theta)$ in equation 2.2). Nonetheless, such diminished utilities may still be higher than the utility under non-strategic policy $\widehat{\theta}^{*}$. This is illustrated empirically in Sec. 5 and Appendix C.
Adjusted preference as a regularizer to promote fairness. We have shown that adjusting weights $k_{1},k_{2},k_{3}$ in learning objective (equation 5) can control the individual behavior and algorithmic fairness. Indeed, we can view this adjustment mechanism as a regularization method: by adjusting weights, we are essentially changing the objective $U(\theta)$ by adding a regularizer, i.e.,
@@ B.5 Estimate Model Parameters
Robustness of results when $q,\epsilon$ are noisy. We also present an experiment to relax the assumption that the decision-maker knows $q,\epsilon$ exactly on FICO data. Instead, they only know $q+\delta$ or $\epsilon+\delta$ where $\delta$ is a Gaussian noise. Table 3 show the results with noisy $q$, while Table 4 show the results with noisy $\epsilon$). The results show adjusting $k$ still works under noisy $q$ and $\epsilon$ although inconsistency exists.
@@ C.1 Additional results on FICO score
Figure 10: Illustration of equation C.2
Experiments with demographic parity as a new fairness metric.
We also reconducted the above experiments with demographic parity (DP) as the new fairness metric. As illustrated in Sec. 4.3, $P_{s}^{\texttt{DP}}(x)=P_{X|S}(x|s)$. Similar to Fig. 4 and Fig. 4, we produce Fig. 11 based on DP, which demonstrate the same patterns as the figures based on Eqopt.
@@ C.2 Results for Gaussian Data
To illustrate the complex nature under different permutations of parameters, with the pre-determined parameters in equation C.2 and $\alpha=0.6$, we vary $q$ and plot both non-strategic optimal thresholds and regular strategic ones with respect to different $\epsilon$ as shown in the bottom plot of Fig. 12, where the lower graphs illustrate Thm. 4.1, i.e. the red line is always under the blue line.
Figure 12: (Non)-strategic optimal threshold. We regularize $y$-axis to be [-5,5] to prevent numerical issues.
We also demonstrate the strategic utility under different combinations of $q,\epsilon$ with pre-determined parameters in equation C.2 and $\alpha=0.3$ or $\alpha=0.6$. Fig. 14 and 14 suggest the complicated nature of regular strategic utility under different parameter combinations. It is possible to have 0,1 or 2 extreme points.
@@
To illustrate 4.5, we demonstrate the effects of adjusting each of $k_{1},k_{2},k_{3}$. According to Fig. 15 and 16, we can see when $k_{1}$ is large enough, the optimal strategic threshold is definitely lower than the optimal non-strategic ones. However, when $\alpha$ is small, we need larger $k_{1}$ to pull $\theta^{*}$ downward. According to Fig. 17 and 18, we can see when the population is majority qualified, adjusting $k_{2}$ is not guaranteed to shift $\theta^{*}$ upward (Fig. 18).
@@
Figure 16: Strategic optimal threshold $\theta^{*}(k_{1})$ after increasing $k_{1}$ while keeping $k_{2},k_{3}$ fixed. Left figure has $q=0.01$ and right figure has $q=0.99$, while both figures have $\alpha=0.3$.We regularize $y$-axis to be [-5,5] to prevent numerical issues.
Figure 17: Strategic optimal threshold $\theta^{*}(k_{2})$ after increasing $k_{2}$ while keeping $k_{1},k_{3}$ fixed. Left figure has $q=0.01$ and right figure has $q=0.99$, while both figures have $\alpha=0.3$. We regularize $y$-axis to be [-5,5] to prevent numerical issues.
@@
We first show a parameter setting satisfying condition 1.(i) in Thm. 4.5. With pre-determined parameters in equation C.2 , we set $q=\epsilon=0.5$ and $\alpha_{a}=0.2,\alpha_{b}=0.25$. This matches the notation tradition in Sec. 4.3 where group $a$ is the disadvantaged group with a lower qualified percentage. Also, because $q+\epsilon\geq 1$, the setting satisfies condition 1.(i) in Thm. 4.5. We first illustrate the manipulation probability under optimal original strategic threshold $\theta_{s}^{*}$ as in Fig. 10. From Fig. 20, we can set $k_{1a}=k_{2a}=1.25$ to let the strategic utility still be larger than the one under non-strategic optimal threshold(i.e. the solid blue line is above the solid red line), while lower the cumulative density dramatically (i.e. the dotted green line) to admit more qualified individuals and disincentivize manipulation (i.e. the dashed blue line). The details of comparisons are shown in Table 7.
@@
With pre-determined parameters in equation C.2 , we set $q=\epsilon=0.25$ and $\alpha_{a}=0.4,\alpha_{b}=0.6$. This matches the notation tradition in Sec. 4.3 where group $a$ is the disadvantaged group with a lower qualified percentage. We first illustrate the manipulation probability under optimal original strategic threshold $\theta_{s}^{*}$ as in Fig. 19. Fig. 19 reveals that 1.(ii) in 4.5 is satisfied because the orange and green points are both located before the extreme large point of $P_{M}(\theta)$. Thus, we could increase $k_{1s}$ to disincentivize manipulation while improving fairness as shown in Fig. 20. From Fig. 20, we can set $k_{1a}=k_{2a}=1.25$ to let the strategic utility still be larger than the one under non-strategic optimal threshold(i.e. the solid blue line is above the solid red line), while lower the cumulative density dramatically (i.e. the dotted green line) to admit more qualified individuals and disincentivize manipulation (i.e. the dashed blue line). In Table 5, We summarize the comparison between non-strategic $\widehat{\theta}^{*}$, original strategic ${\theta}^{*}$, and adjusted strategic ${\theta}^{*}(k_{1})$ (when $k_{1,c}=k_{1,aa}=1.25$). It shows that decision-makers by adjusting preferences can significantly mitigate unfairness and disincentivize manipulation, with only slight decreases in utilities.
@@
Besides, we also show one more parameter setting satisfying condition 2 in Thm. 4.5. With pre-determined parameters in equation C.2 , we also set $q=\epsilon=0.2$ and $\alpha_{a}=0.3,\alpha_{b}=0.35$. This matches the notation tradition in Sec. 4.3 where group $a$ is the disadvantaged group with a lower qualified percentage. Also, based on Fig. 19, $q+\epsilon<1$ and $\alpha_{a},\alpha_{b}<0.5$, the setting satisfies condition 2 in Thm. 4.5. We first illustrates the manipulation probability under optimal original strategic threshold $\theta_{s}^{*}$ and non-strategic threshold $\widehat{\theta_{s}^{*}}$ as in Fig. 19. As shown in Fig. 20, for both groups, we demonstrate the manipulation probability for $\widehat{\theta}^{*},\theta^{*}$ and $\theta(k_{1})$ when $k_{1}$ varies, (non)-strategic utility and cumulative density conditioned on $Y=1$ (i.e. this measures the unfairness based on equal opportunity). This plot suggests we can find suitable $k_{2a}$ and $k_{2b}$ to disincentivize manipulation and promote fairness, while also making the utility higher than the one under non-strategic optimal threshold. From Fig. 20, we can set both $k_{2a}$ and $k_{2b}$ at $1.25$ to let the strategic utility still be larger than the utility under non-strategic optimal threshold (i.e. the solid blue line is above the solid red line), while keeping the cumulative density function closer (i.e. the green dotted line) to mitigate unfairness, and also disincentivize manipulation (i.e. the blue dashed line). The details of comparisons are shown in Table 7.
@@ Appendix D Derivations and Proofs
D.1 Derivations of equation 1
$U_{M}(\theta)$ is the expected utility gain of an unqualified agent if choosing to manipulate: i. If the manipulation is not exposed, the probability of admission is $1-F_{X|Y}(\theta|1)$ because the manipulation leads the agents to get his/her new feature from $P_{X|Y=1}$, which happens at a probability $1-\epsilon$; ii. If the manipulation is exposed, the probability of admission is 0, which happens at a probability $\epsilon$; iii. If the agent does not manipulate, the probability of admission is $1-F_{X|Y}(\theta|0)$ because now his/her feature is from the unqualified population, and keep in mind that the agents will never know the exact values of his/her feature when he/she makes decisions; Then according to the total probability theorem, the expectation of utility gain $U_{M}(\theta)=(1-\epsilon)\cdot(1-F_{X|Y}(\theta|1))+\epsilon\cdot 0-(1-F_{X|Y}(\theta|0))-C_{M}$.
$U_{I}(\theta)$ is the expected utility gain of an unqualified agent if choosing to improve: i. If the improvement succeeds, the probability of admission is $1-F_{X|Y}(\theta|1)$ because the improvement leads the agents to get his/her new feature from $P_{X|Y=1}$, which happens at a probability $q$; ii. If the manipulation is exposed, the probability of admission is $1-F^{I}(\theta)$, which happens at a probability $1-q$; iii. If the agent does not manipulate, the probability of admission is $1-F_{X|Y}(\theta|0)$.
Then according to the total probability theorem, we can derive $U_{I}(\theta)$ as well. Finally, substitute above two terms into $P_{M}(\theta)=\Pr\left(U_{M}(\theta)>U_{I}(\theta)\right)$ and we get equation 1.
@@ D.1 Derivations of equation 1
Assumption 2.2 ensures that $P_{C_{M}-C_{I}}>0$ when $\theta$ in its domain. Thus, we can directly take the derivative inside equation 1, we can get $(1-q)\cdot P^{I}(\theta)-(1-q-\epsilon)P_{X|Y}(\theta|1)$. To get its sign, we only need to consider $(1-q)-(1-q-\epsilon)\frac{P_{X|Y}(\theta|1)}{P^{I}(\theta)}$.